OpenStudy (anonymous):

simplify completely radical 2 and radical 60 over radical 5 ?

OpenStudy (solomonzelman):

\[\huge\color{blue}{\sqrt{2}\frac{\sqrt{60}}{\sqrt{5}}}~~this?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the 2 is next to the 60

OpenStudy (solomonzelman):

\[\huge\color{green}{2\frac{\sqrt{60}}{\sqrt{5}}}\]

OpenStudy (solomonzelman):

yes?

OpenStudy (anonymous):